Using a 0.30 M phosphate buffer with a pH of 8.0, you add 0.70 mL of 0.55 M HCl to 52 mL of the buffer. What is the new pH of the solution? (Enter your answer to three significant figures.)
WebAssign will check your answer for the correct number of significant figures.  

Using a 0.30 M phosphate buffer with a pH of 8.0, you add 0.70 mL of 0.55 M NaOH to 52 mL of the buffer. What is the new pH of the solution? (Enter your answer to three significant figures.)
WebAssign will check your answer for the correct number of significant figures.
Answer #1

Q1. 0.55 M HCl

new pH = 7.92

Q2. 0.55 M NaOH

new pH = 8.10

Explanation

Q1. Total phosphate concentration = 0.30 M

[H2PO4-] + [HPO42-] = 0.30 M ...(1)

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([HPO42-] / [H2PO4-])

8.0 = 7.21 + log([HPO42-] / [H2PO4-])

log([HPO42-] / [H2PO4-]) = 8.0 - 7.21

log([HPO42-] / [H2PO4-]) = 0.79

[HPO42-] / [H2PO4-] = 100.79

[HPO42-] / [H2PO4-] = 6.2

[HPO42-] = 6.2 * [H2PO4-]   ...(2)

Solving equations (1) and (2), we get

[H2PO4-] = 0.0417 M

[HPO42-] = 0.2583 M

moles H2PO4- = (concentration H2PO4-) * (volume buffer)

moles H2PO4- = (0.0417 M) * (52 mL)

moles H2PO4- = 2.17 mmol

Similarly, moles HPO42- = 13.43 mmol

moles HCl added = (concentration HCl) * (volume HCl)

moles HCl added = (0.55 M) * (0.70 mL)

moles HCl added = 0.385 mmol

new moles H2PO4- = (initial moles H2PO4-) + (moles HCl added)

new moles H2PO4- = (2.17 mmol) + (0.385 mmol)

new moles H2PO4- = 2.555 mmol

new moles HPO42- = (initial moles HPO42-) - (moles HCl added)

new moles HPO42- = (13.43 mmol) - (0.385 mmol)

new moles HPO42- = 13.045 mmol

According to Henderson - Hasselbalch equation,

pH = pKa + log([HPO42-] / [H2PO4-])

pH = 7.21 + log(13.045 mmol / 2.555 mmol)

pH = 7.92